OpenStudy (anonymous):

solve y-y1=m(x-x1) for x1.

OpenStudy (anonymous):

Distribute

OpenStudy (anonymous):

y - y1 = mx - mx1 mx1 = mx - y + y1 x1 = (mx - y + y1)/m

OpenStudy (ajprincess):

!.Divide both sides by m 2.Subtract x both sides. 3.Divide by -1 both sides

OpenStudy (anonymous):

How to solve this problem " passes through (-4,2) parrallel to the line whose equation is y=1/2x+5 ??

OpenStudy (ajprincess):

For parallel lines the gradients are equal. So the gradient of the required line=1/2 \(y-y_1=m(x-x_1)\) \(y-2=\frac{1}{2}(x-(-4)\) \(y-2=\frac{1}{2}(x+4)\) solve for y. Can u do?

OpenStudy (anonymous):

No

OpenStudy (ajprincess):

\(y-2=\frac{1}{2}(x+4)\) \(y-2=\large\frac{x}{2}+\frac{4}{2}\) \(y-2=\large\frac{x}{2}+2\) \(y=\large\frac{x}{2}+2+2\) \(y=\large\frac{x}{2}+4\) Is it clear?

OpenStudy (anonymous):

Why does the 1 in the fraction turn to x?

OpenStudy (ajprincess):

When u multiply 1 by x u get x. Is that clear?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yes

OpenStudy (ajprincess):

ok 5yn:)

OpenStudy (anonymous):

What If its passes through (1,-1) parrallel to the line that passes through (4,1) and (2,-3)

OpenStudy (ajprincess):

plz post a question per post.

OpenStudy (ajprincess):

1.Find the gradient \(m_1\)of the line passing through (4,1) and (2,-3) usin the formula \[m_1=\frac{ y_2-y_1 }{ x_2-x_1 }\] \[=\frac{ -3-1 }{ 2-4 }\] 2.For parallel lines gradients are equal. \(m=m_1\) so \[m_1=\frac{ -3-1 }{ 2-4 }\] \[\frac{ -3-1 }{ 2-4 }=\frac{ y-y_1 }{ x-x_1}\] \[\frac{ -3-1 }{ 2-4 }=\frac{ y-(-1) }{ x-1}\]

OpenStudy (anonymous):

Then it turns to 12/2 right??

OpenStudy (ajprincess):

\[\frac{ -3-1 }{ 2-4}=\frac{ y+1 }{ x-1 }\] \[\frac{ -4 }{ -2 }=\frac{ y+1 }{ x-1}\] \[2=\frac{ y+1 }{ x-1}\]

OpenStudy (anonymous):

Oh ok:)